OpenStudy (anonymous):

Two lines, A and B, are represented by the equations given below: Line A: 2x + 2y = 8 Line B: x + y = 4 Which statement is true about the solution to the set of equations? It is (8, 4). It is (4, 8). There is no solution. There are infinitely many solutions.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

Someone helped me with this before and said it was infinite, but I don't think it's the right answer because when I plug in a number it doesn't work

OpenStudy (skullpatrol):

What number are you plugging in?

OpenStudy (anonymous):

I used 2

OpenStudy (anonymous):

wait.

OpenStudy (anonymous):

2 works but any other number doesn't

OpenStudy (skullpatrol):

You can not just use one number, you need two numbers x and y, right?

OpenStudy (anonymous):

yes

OpenStudy (skullpatrol):

Multiply line B by 2 on both sides, what do get?

OpenStudy (anonymous):

Im not sure what you mean

OpenStudy (puppylover52):

its 2 times\[its 2\times2+2\times2=8\]

OpenStudy (skullpatrol):

Line B: x + y = 4 multiplying both sides by 2 2*(x + y) = 2*4 2x + 2y = 8 This^ is exactly the same as line A Therefore they are the same line.

OpenStudy (skullpatrol):

Two lines, A and B, are represented by the equations given below: Line A: 2x + 2y = 8 Line B: x + y = 4 Which statement is true about the solution to the set of equations? It is (8, 4). It is (4, 8). There is no solution. There are infinitely many solutions. Answer

OpenStudy (skullpatrol):

All the points that satisfy line A will also satisfy line B because they are the same line. Recall to satisfy an equation means those values of the variable that make it a true statement.